OpenStudy (anonymous):

What is (-m)^-3n If m=3 and n=-25?

OpenStudy (shane_b):

First, welcome to OpenStudy! Is this the problem?\[(-m)^{-3n}\] or is it \[(-m)^{-3}n\]?

jhonyy9 (jhonyy9):

so i think this wann being (-m)^-3 *n so than this mean that (-m)^-3 = 1/ (-m)^3 *n = n/(-m)^3 so if m=3 and n=-25 than this will be - 25 - 25 25 ------ = ------ = ---- (-3)^3 - 27 27 hope you will understand it and that this will help you good luck bye